Find the probability that the sum is as stated when a pair of dice is rolled. (Enter your answers as fractions.)
A. odd and less than 5
B. odd or less than 5

Answer by mananth(16949)


When two dice are rolled total possible outcomes is 3
A. odd and less than 5
Sum odd on the top face
(1,2),(1,4),(1,6),
(2,1),(2,3),(2,5),
(3,2),(3,4),(3,6)
(4,1),(4,3),(4,5)
(5,2),(5,4),(5,6)
(6,1),(6,3),(6,5)
odd and less than 5
(1,2)(2,1)
P = 2/36 =1/18---------A

B. odd or less than 5
Choose from above all odd and less than 5